# Astrometry-Only Detection of Microlensing Events with Gaia

### v1

@@ 3.1 Parameter recovery accuracy and degeneracies
We applied GAME Filter to the lens_G14_N281_DR4 dataset ($G_{0}=14$, $N_{\mathrm{v}}=281$, and DR4 observational run) to estimate the accuracy of parameter recovery and investigate degeneracies between parameters. We found that the percentages of recovered events are: $P_{\mathrm{rec}}=61\%$, $P_{20}=55\%$, and $P_{10}=39\%$ (see Table 3).
Figure 2: 2D histograms of values of the parameters obtained after minimization as a function of their true values for the lens_G14_N281_DR4 dataset (see Tables 1 and 3). The colourbar corresponds to the number of events in individual bins.
@@
Figure 3: Corner plots for ratios $\mathcal{R}$ between true and minimized values of individual parameters for recovered events in the dataset lens_G14_N281_DR4 (see Tables 1 and 3).
@@ 3.2 Parameter recovery for sources with different magnitudes
We tested GAME Filter on three datasets, namely lens_G14_N281_DR4, lens_G16.5_N281_DR4, and lens_G19_N281_DR4, which correspond to $G_{0}=14$, 16.5, and 19, respectively. The results of these tests are summarised in Table 3. As the magnitude increases, noise increases, and the source becomes fainter, the observed signal becomes weaker. Consequently, the number of recovered events decreases and the error of parameter estimation increases, indicated by progressively lower values of $P_{\mathrm{rec}}$, $P_{20}$, and $P_{10}$ as $G_{0}$ increases.
@@
Figure 4: Distributions of parameter values after minimization for recovered events for datasets lens_G14_N281_DR4, lens_G16.5_N281_DR4, and lens_G19_N281_DR4 corresponding to $G_{0}=$ 14 (blue), 16.5 (green), and 19 (red), respectively. The values in individual bins denote the number of recovered events within the corresponding bin range. The integral of distributions is normalized to $P_{\mathrm{rec}}$ (see Table 3).
Figure 5: Relative errors for recovered events for datasets lens_G14_N281_DR4, lens_G16.5_N281_DR4, and lens_G19_N281_DR4 corresponding to $G_{0}=$ 14 (blue), 16.5 (green), and 19 (red). The errors are calculated as the difference between minimized and true parameter values divided by the minimized parameter values. The error values are first binned and then calculated as the mean in each bin.
@@ 3.3 Parameter recovery for events with different numbers of Gaia visits
We tested GAME Filter on datasets lens_G14_N281_DR4, lens_G14_N209_DR4, and lens_G14_N91_DR4, which correspond to regions with different numbers of Gaia visits $N_{\mathrm{v}}=281$, 209, and 91. The results are summarised in Table 3, showing that the changes in $N_{\mathrm{v}}$ do not significantly affect $P_{\mathrm{rec}}$. Additionally, $P_{20}$ and $P_{10}$ remain unaffected when $N_{\mathrm{v}}$ decreases to at least 209. When $N_{\mathrm{v}}$ is further reduced to 91, $P_{20}$ and $P_{10}$ show only a $\approx 5$% change.
@@
We also calculated the relative errors for individual parameters after minimisation and find that the errors increase slightly with decreasing $N_{\mathrm{v}}$. Since the impact is not significant, this indicates that GAME Filter performs robustly even with fewer observations. However, a higher number of observations generally improves the accuracy of parameter recovery.
Figure 6: Distributions of parameter values after minimization for recovered events for datasets lens_G14_N281_DR4, lens_G14_N209_DR4, and lens_G14_N91_DR4 corresponding to $N_{\mathrm{v}}=281$ (blue), 209 (green), and 91 (red). The values in individual bins denote the number of recovered events within the corresponding bin range. The integral of distributions is normalized to $P_{\mathrm{rec}}$ (see Table 3).
@@ 3.4 Parameter recovery for events with the maximum microlensing signal beyond Gaia observational run
We tested GAME Filter on datasets lens_G14_N281_DR4 and lens_G14_N281_extend, which correspond to datasets within Gaia DR4 with $t_{0}\in[2014.5,2020]\,$yr and datasets extending beyond DR4 with $t_{0}\in[2010,2024]\,$yr, respectively. The results of these tests are summarised in Table 3. We see that increasing the $t_{0}$ range decreases $P_{\mathrm{rec}}$, $P_{20}$ and $P_{10}$ by $\approx 15\%$, $\approx 20\%$, and $15\%$, respectively.
In Figure 7, we show histograms of the parameter values after minimisation for recovered events for different ranges of $t_{0}$. As $t_{0}$ extends beyond the DR4 run, $P_{\mathrm{rec}}$ decreases. This is because a significant part of the microlensing event can fall outside the Gaia observation time span, reducing the amount of available data and making parameter recovery more challenging. This effect is most apparent for events with a short $t_{\mathrm{E}}$. Furthermore, the histogram of $u_{0}$ transitions from a bimodal distribution to a single peak as the $t_{0}$ range increases. There are also fewer recovered events for large $\theta_{\mathrm{E}}$, as larger $\theta_{\mathrm{E}}$ corresponds to longer lens paths and event durations, which means that a smaller portion of the event is observed if $t_{0}$ is outside the DR4 time span. We also see fewer recovered events for microlensing parallaxes close to -1 and 1, since these events have more variations, or ”wiggles” in the observed astrometric signal, which makes the parameter recovery more challenging. We discuss the physical mechanism for these differences in Section 4.1.
Figure 7: Distributions of parameter values after minimization for recovered events for datasets lens_G14_N281_DR4 and lens_G14_N209_extended, corresponding to $t_{0}\in[2014.5,2020]\,$yr and $t_{0}\in[2010,2024]\,$yr, respectively. The values in individual bins denote the number of recovered events within the corresponding bin range. The integral of distributions is normalized to $P_{\mathrm{rec}}$ (see Table 3).
@@
Figure 8: Relative errors for recovered events for datasets lens_G14_N281_DR4 and lens_G14_N209_extended, corresponding to $t_{0}\in[2014.5,2020]\,$yr and $t_{0}\in[2010,2024]\,$yr, respectively. The errors are calculated as the difference between minimized and true parameter values divided by the minimized parameter values. The error values are first binned and then calculated as the mean in each bin.
@@ 3.5 Parameter recovery for the false positive detections
To evaluate the performance of GAME Filter in distinguishing microlensing events from other astrometric phenomena, we tested it on the binary_G14_N281_DR4 dataset, which corresponds to binary system events. The results of these tests are summarised in Table 3. We determined that the percentage of false positive identifications is very low, with $P_{\mathrm{rec}}=5\%$ and $P_{20}=P_{10}=0\%$.
@@
Figure 9: Distributions of parameter values after minimization for binary system events in the binary_G14_N281_DR4 dataset (see Tables 2 and 3). The values in individual bins denote the number of recovered events within the corresponding bin range. The integral of distributions is normalized to $P_{\mathrm{rec}}$ (see Table 3).
@@ 4.1 Parameter recovery accuracy
The microlensing ellipse is also affected by $\pi_{\mathrm{EE}}$ and $\pi_{\mathrm{EN}}$, which determine both the orientation of the ellipse in the sky and the parallax of the lens $\pi_{\mathrm{L}}$ (see Equation 5). Higher values of $\pi_{\mathrm{EE}}$ and $\pi_{\mathrm{EN}}$ lead to a larger $\pi_{\mathrm{E}}$ and consequently to a higher $\pi_{\mathrm{L}}$, assuming a fixed $\varpi$. When the lens is closer to the observer, its parallactic motion introduces more variations, or ”wiggles,” in the observed astrometric signal. This added complexity makes parameter estimation more challenging, requiring more measurements to accurately constrain the parameters. This effect is particularly pronounced when $t_{0}$ lies outside of the Gaia observational window, as shown in Figure 7.
@@
iv) In Gaia DR4, different regions in the sky are observed on average $\approx 100$ times. We find that for $N_{\mathrm{v}}\geq 91$, a decrease in $N_{\mathrm{v}}$ leads to only a slight increase in relative errors. For example, when $N_{\mathrm{v}}$ decreases from 281 to 91, the percentage of parameters with their estimated values with 10% of their true values decreases by $\approx 5\%$ (see Table 3).
v) Gaia observes various types of astrophysical astrometric events, including single stars, binary star systems, microlensing events, and other transient phenomena. Among these, astrometric binary systems are the primary source of potential false positives in microlensing detection. We find that fewer than 5% of binary events are incorrectly identified as weak microlensing events. These false positives typically mimic the astrometric signature of weak microlensing events with low $\theta_{\mathrm{E}}$ and high $|u_{0}|$.
@@ 4.2 Constraining the lens properties
To evaluate the sensitivity of the GAME Filter to detect lenses in the Milky Way with varying distances $D_{\mathrm{L}}$ and masses $M_{\mathrm{L}}$, we applied the GAME Filter to a new dataset of $N_{\mathrm{all}}=50,000$ microlensing events. For each event, $M_{\mathrm{L}}$ is randomly selected from a logarithmically spaced interval between $0.1$ and $20\,\mathrm{M_{\odot}}$, and $D_{\mathrm{L}}$ is uniformly sampled from $0.1$ to $7.9\,\mathrm{kpc}$. The angle between the microlensing parallax components $\pi_{\mathrm{EE}}$ and $\pi_{\mathrm{EN}}$ is selected from a uniform distribution between 00 and $2\pi$. Additional parameters are sampled uniformly within their respective ranges (see Table 1) with the exception of $u_{0}$ where we changed the range to -3 and 3. Using these values, the microlensing parallax $\pi_{\mathrm{E}}$ and the angular Einstein radius $\theta_{\mathrm{E}}$ are calculated from $\pi_{\mathrm{S}}$, $M_{\mathrm{L}}$, and $D_{\mathrm{L}}$ (see Equations 1 and 5). GAME Filter recovers $N_{20\%}=11,456$ events with parameter values minimised within 20% of their true values, corresponding to approximately $22\%$ of $N_{\mathrm{all}}$.
@@ 4.3 Caveats
The optimisation process in GAME Filter may occasionally converge to a local minimum instead of the global minimum, potentially leading to incorrect parameter estimates. To mitigate this, we used multiple initial guesses for parameters $u_{0}$, $\pi_{\mathrm{EE}}$, and $\pi_{\mathrm{EN}}$ for which we lack initial estimates (see Appendix A). Specifically, we used 10 different values for $u_{0}$ and 5 different values for both $\pi_{\mathrm{EE}}$ and $\pi_{\mathrm{EN}}$ within a defined interval. Although this approach improves parameter recovery accuracy, it does not cover the entire parameter space. To evaluate this, we further tested GAME Filter on a subset of lens_G14_N281_DR4 dataset with 5000 events, increasing the number of initial guesses for $u_{0}$, $\pi_{\mathrm{EE}}$ and $\pi_{\mathrm{EN}}$ to 13, 7, and 7, respectively. We also applied this approach to the parameters $\theta_{\mathrm{E}}$, $t_{0}$, and $t_{\mathrm{E}}$, where we considered 5 different values for individual parameters. We found no significant improvement compared to our original method. Due to this, we conclude that our original approach strikes a balance between computational efficiency, thorough parameter space exploration, and accurate parameter estimation.
@@
and similarly for the proper motions. When $f_{\mathrm{bl}}$ is near 1, the observed signal mainly reflects the motion of the source, while values close to 0 reflect the motion of the lens. Intermediate values represent mixed cases where accurate estimation of the lens parallax requires additional source distance measurements, e.g. spectroscopic typing. In our study, we considered $f_{\mathrm{bl}}=1$ to determine the lower boundary for the Einstein radius $\theta_{E}$. This, in turn, establishes a minimum estimate for both the distance and mass of the lens. If $f_{\mathrm{bl}}$ were less than 1, indicating a significant contribution of light from the lens, the actual distance and mass would likely be higher than the values estimated under this assumption.
@@ 5 Conclusions
  • Values of MUWE and L2 optimality error after minimization can be used to filter recovered events where GAME Filter reached the desired convergence.
@@
  • Binary systems represent a potential contaminant for GAME Filter’s performance. However, we find that only $\approx 5$% of binary events get falsely identified as weak microlensing events.
The development of GAME Filter demonstrates the astrometric potential in microlensing events. The work presented here suggests that astrometric data can not only complement photometric detections but also enable microlensing detections and the determination of lens parameters independently, even in the absence of photometric signals. Therefore, this study paves the way for future astrometric investigations that fully leverage the capabilities of Gaia data.

### v2

@@ 3.1 Parameter recovery accuracy and degeneracies
We applied GAME Filter to the lens_G14_N281_DR4 mock dataset ($G_{0}=14$, $N_{\mathrm{v}}=281$, and DR4 observational run) to estimate the accuracy of parameter recovery and investigate degeneracies between parameters. We found that the percentages of recovered events are: $P_{\mathrm{rec}}=61\%$, $P_{20}=55\%$, and $P_{10}=39\%$ (see Table 3).
Figure 2: 2D histograms of values of the parameters obtained after minimisation as a function of their true values for the lens_G14_N281_DR4 mock dataset (see Tables 1 and 3). The colourbar corresponds to the number of events in individual bins.
@@
Figure 3: Corner plots for ratios $\mathcal{R}$ between true and minimised values of individual parameters for recovered events in the mock dataset lens_G14_N281_DR4 (see Tables 1 and 3).
@@ 3.2 Parameter recovery for sources with different magnitudes
We tested GAME Filter on three mock datasets, namely lens_G14_N281_DR4, lens_G16.5_N281_DR4, and lens_G19_N281_DR4, which correspond to $G_{0}=14$, 16.5, and 19, respectively. The results of these tests are summarised in Table 3. As the magnitude increases, noise increases, and the source becomes fainter, the observed signal becomes weaker. Consequently, the number of recovered events decreases and the error of parameter estimation increases, indicated by progressively lower values of $P_{\mathrm{rec}}$, $P_{20}$, and $P_{10}$ as $G_{0}$ increases.
@@
Figure 4: Distributions of parameter values after minimisation for recovered events for mock datasets lens_G14_N281_DR4, lens_G16.5_N281_DR4, and lens_G19_N281_DR4 corresponding to $G_{0}=$ 14 (blue), 16.5 (green), and 19 (red), respectively. The values in individual bins denote the number of recovered events within the corresponding bin range. The integral of distributions is normalized to $P_{\mathrm{rec}}$ (see Table 3).
Figure 5: Relative errors for recovered events for mock datasets lens_G14_N281_DR4, lens_G16.5_N281_DR4, and lens_G19_N281_DR4 corresponding to $G_{0}=$ 14 (blue), 16.5 (green), and 19 (red). The errors are calculated as the difference between the minimised and true parameter values divided by the minimised parameter values. The error values are first binned and then calculated as the mean in each bin.
@@ 3.3 Parameter recovery for events with different numbers of Gaia visits
We tested GAME Filter on mock datasets lens_G14_N281_DR4, lens_G14_N209_DR4, and lens_G14_N91_DR4, which correspond to regions with different numbers of Gaia visits $N_{\mathrm{v}}=281$, 209, and 91. The results are summarised in Table 3, showing that the changes in $N_{\mathrm{v}}$ do not significantly affect $P_{\mathrm{rec}}$. Additionally, $P_{20}$ and $P_{10}$ remain unaffected when $N_{\mathrm{v}}$ decreases to at least 209. When $N_{\mathrm{v}}$ is further reduced to 91, $P_{20}$ and $P_{10}$ show only a $\approx 5$% change.
@@
We also calculated the relative errors for individual parameters after minimisation and found that the errors increase slightly with decreasing $N_{\mathrm{v}}$. Since the impact is not significant, this indicates that GAME Filter performs robustly even with fewer observations. However, a higher number of observations generally improves the accuracy of parameter recovery.
Figure 6: Distributions of parameter values after minimisation for recovered events for mock datasets lens_G14_N281_DR4, lens_G14_N209_DR4, and lens_G14_N91_DR4 corresponding to $N_{\mathrm{v}}=281$ (blue), 209 (green), and 91 (red). The values in individual bins denote the number of recovered events within the corresponding bin range. The integral of distributions is normalized to $P_{\mathrm{rec}}$ (see Table 3).
@@ 3.4 Parameter recovery for events with the maximum microlensing signal beyond Gaia observational run
We tested GAME Filter on mock datasets lens_G14_N281_DR4 and lens_G14_N281_extend, which correspond to mock dataset within Gaia DR4 with $t_{0}\in[2014.5,2020]\,$yr and mock dataset extending beyond DR4 with $t_{0}\in[2010,2024]\,$yr, respectively. The results of these tests are summarised in Table 3. We see that increasing the $t_{0}$ range decreases $P_{\mathrm{rec}}$, $P_{20}$ and $P_{10}$ by $\approx 15\%$, $\approx 20\%$, and $15\%$, respectively.
In Figure 7, we show histograms of the parameter values after minimisation for recovered events for different ranges of $t_{0}$. As $t_{0}$ extends beyond the DR4 run, $P_{\mathrm{rec}}$ decreases. This is because a significant part of the microlensing event can fall outside the Gaia observation time span, reducing the amount of available data and making parameter recovery more challenging. This effect is most apparent for events with a short $t_{\mathrm{E}}$. Furthermore, the histogram of $u_{0}$ transitions from a bimodal distribution to a single peak as the $t_{0}$ range increases. There are also fewer recovered events for large $\theta_{\mathrm{E}}$, as larger $\theta_{\mathrm{E}}$ corresponds to longer lens paths and event durations, which means that a smaller portion of the event is observed if $t_{0}$ is outside the DR4 time span. We also see fewer recovered events for microlensing parallaxes close to -1 and 1, since these events have more variations, or “wiggles” in the observed astrometric signal, which makes the parameter recovery more challenging. We discuss the physical mechanism for these differences in Section 4.1.
Figure 7: Distributions of parameter values after minimisation for recovered events for mock datasets lens_G14_N281_DR4 and lens_G14_N209_extended, corresponding to $t_{0}\in[2014.5,2020]\,$yr and $t_{0}\in[2010,2024]\,$yr, respectively. The values in individual bins denote the number of recovered events within the corresponding bin range. The integral of distributions is normalized to $P_{\mathrm{rec}}$ (see Table 3).
@@
Figure 8: Relative errors for recovered events for mock datasets lens_G14_N281_DR4 and lens_G14_N209_extended, corresponding to $t_{0}\in[2014.5,2020]\,$yr and $t_{0}\in[2010,2024]\,$yr, respectively. The errors are calculated as the difference between minimised and true parameter values divided by the minimised parameter values. The error values are first binned and then calculated as the mean in each bin.
@@ 3.5 Parameter recovery for the false positive detections
To evaluate the performance of GAME Filter in distinguishing microlensing events from other astrometric phenomena, we tested it on the binary_G14_N281_DR4 mock dataset, which corresponds to binary system events. The results of these tests are summarised in Table 3. We determined that the percentage of false positive identifications is very low, with $P_{\mathrm{rec}}=5\%$ and $P_{20}=P_{10}=0\%$.
@@
Figure 9: Distributions of parameter values after minimisation for binary system events in the binary_G14_N281_DR4 mock dataset (see Tables 2 and 3). The values in individual bins denote the number of recovered events within the corresponding bin range. The integral of distributions is normalized to $P_{\mathrm{rec}}$ (see Table 3).
@@ 4.1 Parameter recovery accuracy
The microlensing ellipse is also affected by $\pi_{\mathrm{EE}}$ and $\pi_{\mathrm{EN}}$, which determine both the orientation of the ellipse in the sky and the parallax of the lens $\pi_{\mathrm{L}}$ (see Equation 7). Higher values of $\pi_{\mathrm{EE}}$ and $\pi_{\mathrm{EN}}$ lead to a larger $\pi_{\mathrm{E}}$ and consequently to a higher $\pi_{\mathrm{L}}$, assuming a fixed $\varpi$. When the lens is closer to the observer, its parallactic motion introduces more variations in the observed astrometric signal. This added complexity makes parameter estimation more challenging, requiring more measurements to accurately constrain the parameters. This effect is particularly pronounced when $t_{0}$ lies outside of the Gaia observational window, as shown in Figure 7.
@@
iv) In Gaia DR4, different regions in the sky are observed on average $\approx 100$ times. We find that for $N_{\mathrm{v}}\geq 91$, a decrease in $N_{\mathrm{v}}$ leads to only a slight increase in relative errors. For example, when $N_{\mathrm{v}}$ decreases from 281 to 91, the percentage of parameters with their estimated values within 10% of their true values decreases by $\approx 5\%$ (see Table 3).
[…]
@@ 4.2 Constraining the lens properties
To evaluate the sensitivity of the GAME Filter to detect lenses in the Milky Way with varying distances $D_{\mathrm{L}}$ and masses $M_{\mathrm{L}}$, we applied the GAME Filter to a new mock dataset of $N_{\mathrm{all}}=50,000$ microlensing events. For each event, $M_{\mathrm{L}}$ is randomly selected from a logarithmically spaced interval between $0.1$ and $20\,\mathrm{M_{\odot}}$, and $D_{\mathrm{L}}$ is uniformly sampled from $0.1$ to $7.9\,\mathrm{kpc}$. The angle between the microlensing parallax components $\pi_{\mathrm{EE}}$ and $\pi_{\mathrm{EN}}$ is selected from a uniform distribution between 00 and $2\pi$. Additional parameters are sampled uniformly within their respective ranges (see Table 1) with the exception of $u_{0}$ where we changed the range to -3 and 3. Using these values, the microlensing parallax $\pi_{\mathrm{E}}$ and the angular Einstein radius $\theta_{\mathrm{E}}$ are calculated from $\pi_{\mathrm{S}}$, $M_{\mathrm{L}}$, and $D_{\mathrm{L}}$ (see Equations 1 and 7). GAME Filter recovers $N_{20\%}=11,456$ events with parameter values minimised within 20% of their true values, corresponding to approximately $22\%$ of $N_{\mathrm{all}}$.
@@ 4.3 Caveats
The optimisation process in GAME Filter may occasionally converge to a local minimum instead of the global minimum, potentially leading to incorrect parameter estimates. To mitigate this, we used multiple initial guesses for parameters $u_{0}$, $\pi_{\mathrm{EE}}$, and $\pi_{\mathrm{EN}}$ for which we lack initial estimates (see Appendix A). Specifically, we used 10 different values for $u_{0}$ and 5 different values for both $\pi_{\mathrm{EE}}$ and $\pi_{\mathrm{EN}}$ within a defined interval. Although this approach improves parameter recovery accuracy, it does not cover the entire parameter space. To evaluate this, we further tested GAME Filter on a subset of lens_G14_N281_DR4 mock dataset with 5000 events, increasing the number of initial guesses for $u_{0}$, $\pi_{\mathrm{EE}}$ and $\pi_{\mathrm{EN}}$ to 13, 7, and 7, respectively. We also applied this approach to the parameters $\theta_{\mathrm{E}}$, $t_{0}$, and $t_{\mathrm{E}}$, where we considered 5 different values for individual parameters. We found no significant improvement compared to our original method. Due to this, we conclude that our original approach strikes a balance between computational efficiency, thorough parameter space exploration, and accurate parameter estimation.
@@
and similarly for the proper motions. When $f_{\mathrm{bl}}$ is near 1, the observed signal mainly reflects the motion of the source, while values close to 0 reflect the motion of the lens. Intermediate values represent mixed cases where accurate estimation of the lens parallax requires additional source distance measurements, e.g., spectroscopic typing. In our study, we considered $f_{\mathrm{bl}}=1$ to determine the lower boundary for the Einstein radius $\theta_{E}$. This, in turn, establishes a minimum estimate for both the distance and mass of the lens. If $f_{\mathrm{bl}}$ were less than 1, indicating a significant contribution of light from the lens, the actual distance and mass would likely be higher than the values estimated under this assumption.
@@ 5 Conclusions
  • Values of MUWE and L2 optimality error after minimisation can be used to filter recovered events where GAME Filter reached the desired convergence.
@@
  • Binary systems represent a potential contaminant for GAME Filter’s performance. We found that only $\approx 5$% of binary events get falsely identified as weak microlensing events for $G_{0}=14$. However, we assumed uniform orbital parameter distributions, likely overestimating the confusion rate by underrepresenting short-period and overrepresenting long-period binaries. Nevertheless, we found that the false positives exhibit uniform distributions of parameters, suggesting that statistical noise rather than particular orbital configurations is the dominant cause of confusion.
We presented our results for Gaia DR4, where GAME Filter demonstrates robust performance in detecting astrometric microlensing events. For DR5, we expect a similar overall performance. Additionally, we expect that the longer observing time will allow the detection of longer-duration events, particularly those with larger $\theta_{\mathrm{E}}$, which may have been missed in DR4 due to insufficient sampling. Such events are expected to be less frequent, as lenses with large masses are inherently rarer. However, these events would be of special astrophysical interest, as they are more likely to involve massive black holes. Additionally, we note that a quantitative estimate of the number of astrometric microlensing events detectable in DR4 requires detailed simulations with realistic Galactic population models. Such an estimate is beyond the scope of this paper. We plan to provide detailed predictions of event rates using realistic populations in a future study (Kaczmarek, et al., in prep.).
The work presented here suggests that astrometric data can not only complement photometric detections but also enable microlensing detections and the determination of lens parameters independently, even in the absence of photometric signals. Therefore, this study paves the way for future astrometric investigations that fully leverage the capabilities of Gaia data.